Please show work8.)The Walking Dead is an American post-apocalyptic horror drama TV series about a sheriffs deputy who awakens from a coma
Answer #1

Part a

The sample correlation coefficient between the dependent variable IMBD rating and independent variable number of episodes is given as r = 0.427074909. This means there is a considerable or intermediate positive linear relationship exists between the given two variables IMBD rating and number of episodes.

Part b

Here, we have to use t test for the population correlation coefficient. The null and alternative hypotheses for this test are given as below:

Null hypothesis: H0: There is no any statistically significant correlation between the two variables IMBD ratings and number of episodes.

Alternative hypothesis: Ha: There is a statistically significant correlation between the two variables IMBD ratings and number of episodes.

Part c

We assume level of significance for this test as α = 0.05.

The test statistic formula is given as below:

t = r*sqrt(n – 2)/sqrt(1 – r^2)

We are given

n = 6

r = 0.427074909

t = 0.427074909*sqrt(6 - 2)/sqrt(1 - 0.427074909^2)

t = 0.854149818/0.904216247

t = 0.944630027

df = n – 2 = 6 – 2 = 4

P-value = 0.3983

(by using t-table or excel)

P-value > α = 0.05

So, we do not reject the null hypothesis

There is insufficient evidence to conclude that there is a statistically significant correlation between the two variables IMBD ratings and number of episodes.

Part d

The regression model by using excel is given as below:

SUMMARY OUTPUT

Regression Statistics

Multiple R

0.427074909

R Square

0.182392978

Adjusted R Square

-0.022008778

Standard Error

0.635368765

Observations

6

ANOVA

df

SS

MS

F

Significance F

Regression

1

0.360226131

0.360226131

0.892325885

0.398335369

Residual

4

1.614773869

0.403693467

Total

5

1.975

Coefficients

Standard Error

t Stat

P-value

Lower 95%

Upper 95%

Intercept

8.15879397

0.488662983

16.69615717

7.53997E-05

6.802048023

9.515539917

Episode number

0.013567839

0.014363125

0.944630026

0.398335369

-0.026310589

0.053446267

The required regression equation or line of best fit is given as below:

IMBD rating = 8.15879397 + 0.013567839*Episode number

Where,

Y-intercept = 8.15879397

Slope = 0.013567839

Y-intercept indicates that there is a 8.16 rating when no episode number is considered.

Slope of this regression line indicate that there is an increment of 0.01 approximately in the IMBD rating as there is increment of one in episode number.

Part e

We are given

Episode number = 26

IMBD rating = 8.15879397 + 0.013567839*Episode number

IMBD rating = 8.15879397 + 0.013567839*26

IMBD rating = 8.511557784

IMBD rating = 8.5 approximately

Actual rating = 8.6

Error in prediction = actual rating – predicted rating = 8.6 – 8.5 = 0.1

Error in prediction = 0.1